OpenStudy (anonymous):

What is the probability of rolling a 4 twice in a row with a six-sided die, numbered 1 through 6?

OpenStudy (anonymous):

oh.... ok i think that helps.... thanks!

OpenStudy (anonymous):

i think it might be \(\frac{1}{6}\times \frac{1}{6}\)

OpenStudy (anonymous):

i read it as "rolling a "4" twice in a row"

OpenStudy (anonymous):

yeah its 1/6....

OpenStudy (anonymous):

actually it is \(\frac{1}{36}\)

OpenStudy (anonymous):

With who?

OpenStudy (slaaibak):

It's 1/6 * 1/6, like satellite73 said

OpenStudy (anonymous):

I believe it means rolling a number 4 twice: P ( 4 and 4) = 1/6 * 1/6 = 1/ 36

OpenStudy (slaaibak):

The first problem clearly states either heads twice OR tails twice. Which is something different. There two outcomes are looked at. This problem looks at one outcome, namely 4 rolled twice

OpenStudy (anonymous):

it is 1/36

OpenStudy (callisto):

I also agree that it is 1/36 The first time you row, you need to get a 4 , so P(4) = 1/6 The second time you row, you need to get a 4 , so P(4) = 1/6 Since you need to get a '4' in the first time AND the second time P(4 and 4) = 1/6 x1/6 = 1/36 For the example given by Luis, P(both head) = 1/2 x 1/2 = 1/4 P(both tail) = 1/2 x 1/2 = 1/4 Since it requires either both tails OR both head P(both tails or both heads) = 1/4 +1/4 = 1/2